HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

CONSTRUCTION AND PRELIMINARY RESULTS OF A KU-BAND CHIRPED-PULSE MICROWAVE FOURIER TRANSFORM SPECTROMETER

\underline{S. COLLIGNON} , Institute of Condensed Matter and Nanosciences, Université

catholique de Louvain, Louvain-la-Neuve, B-1348, Belgium; B. M. HAYS, Université de Lille, CNRS, UMR8523—PhLAM—Physique des Lasers Atomes et Molécules, F-59000 Lille, France; D. LEDERER, Institute of Information and Communication Technologies, Electronics and Applied Mathematics, Université catholique de Louvain, Louvain-la-Neuve, B-1348, Belgium; M. DAMAN, C.

LAUZIN, Institute of Condensed Matter and Nanosciences, Université catholique de Louvain, Louvain-la-Neuve, B-1348, Belgium

The chirped-pulse Fourier transform microwave (CP-FTMW) spectroscopy technique is an efficient tool for the rapid measurement of broadband rotational spectra ^{1}. Fast, state-of-the-art arbitrary waveform generators (AWG) and oscilloscopes are usually required for CP-FTMW spectrometers to induce and detect the free induction decay (FID) in a gas of molecules.

Alternatively, the CP-FTMW spectrometer currently developed at UCLouvain makes use of a 2.5 GSa/s direct digital synthesizer in place of an AWG, following the work of Ref. ^{2}. It also leverages heterodyne detection to allow the acquisition of the FIDs with a 1 GSa/s oscilloscope. A quadrature demodulation technique is implemented to discriminate between the lower and upper sidebands of the probed frequency range ^{3}. A computer program was developed to automate the scan of the Ku-band (12−18 GHz) and remove the spurious content from the spectra.

Methanol spectra have been measured in the Ku-band to assess the performances of the spectrometer. The methanol sample was first studied at room temperature in a waveguide cell. The 12−18 GHz frequency range was covered by concatenation of multiple 500 MHz sub-spectra, all acquired and averaged 10^{5} times in ∼ 13 minutes. In this configuration, the acquisition rate, precision and sensitivity of the instrument were evaluated. Based on Ref. ^{4}, we fitted the FID in the time domain and extracted the resonant frequency and the T_{2} coherence time. First measurements performed in a supersonic expansion will also be presented.

1doi:10.1063/1.2919120, G. G. Brown et al., Rev. Sci. Instrum., 79, 053103 (2008).

2doi:10.1063/1.4818137, I. A. Finneran et al., Rev. Sci. Instrum., 84, 083104 (2013).

3doi:10.1039/C1CP22197C, J.-U. Grabow et al., Phys. Chem. Chem. Phys., 13, 21063-21069 (2011).

4doi:10.1016/j.jqsrt.2020.107001, B. M. Hays et al., J. Quant. Spectrosc. Radiat. Transf., 250, 107001 (2020).

p-number: p233  Submitted on Wed, 21 Jun 2023 18:32:16 +0200